An investment portfolio of stocks and bonds, with an initial investment of $120,000 gained $30,800 over its life. If the stocks gained 32% and the bonds gained 16%, how much was invested in each?
Answer by checkley77(12844)

.32x+.16(120,000-x)=30,800
.32x+19,200-.16x=30,800
.16x=30,800-19,200
.16x=11,600
x=11,600/.16
x=72,500 amount that gained 32%.
120,000-72,500=47,500 amount that gained 16%.
Proof:
.32*72,500+.16*47,500=30,800
23,200+7,600=30,800
30,800=30,800